In a particular bag of of skittles, there are 13 red, 12 orange, 10 green, 12 yellow, and 11 purple skittles.
Answer as a fraction or w/o rounding
1. Selecting a yellow skittle _______
2. Selecting a red or purple skittle ______
3. Not selecting a green skittle ______
4. Selecting two consecutive orange skittles _____

Answer by math_tutor2020(3835)


Problem 1

There are 12 yellow out of 13+12+10+12+11 = 58 total.

This forms the fraction 12/58 = (2*6)/(2*29) = 6/29 which is the probality of selecting a yellow skittle.

====================================================
Problem 2

We have 13 red + 11 purple = 24 skittles we want to select out of 58 total

24/58 = (2*12)/(2*29) = 12/29 is the probability of selecting either a red or purple skittle (note the events are mutually exclusive).

====================================================
Problem 3

There are 10 green and 58-10 = 48 non-green skittles.

48/58 = (2*24)/(2*29) = 24/29 is the probability of selecting a skittle that isn't green.

====================================================
Problem 4

This is assuming you aren't putting the first one back, or no replacement has been made.

12/58 represents the probability of selecting an orange skittle (12 orange out of 58 total)

11/57 is the probability of selecting another orange skittle (12-1 = 11 orange left, out of 58-1 = 57 total left)

Multiply the probabilities
(12/58)*(11/57)
(12*11)/(58*57)
132/3306
22/551 is the probability of selecting two orange skittles in a row (no replacement).